OpenStudy (nathanjhw):

The graph of the equation 2x=y^2 from A(0,0) to B(2,2) is revolved about the x-axis.The surface area of the resulting solid (in square units and to three decimal places) is:

OpenStudy (nathanjhw):

18.088 21.322 3.3934 18.585 22.062

OpenStudy (nathanjhw):

@iambatman

OpenStudy (nathanjhw):

@freckles

OpenStudy (freckles):

is that really what is written? like what equation are they talking?

OpenStudy (nathanjhw):

I am sorry I fixed the error. It is = not +

OpenStudy (freckles):

ok do you know the surface area formula ?

OpenStudy (nathanjhw):

no

OpenStudy (freckles):

I'm going to assume your stuff is bounded by 2x=y^2 and y=0 and the points you gave anyway you should know the formulas at least you might want to use these notes in the future http://tutorial.math.lamar.edu/Classes/CalcII/SurfaceArea.aspx but anyways I decided to put this integral in terms of y since it seemed easier since were rotating about the x-axis and your equation x=1/2*y^2 is already solved for x just find your limits now (these are actually already given ) \[\int\limits_{c}^{d}2 \pi y \sqrt{1+(\frac{dx}{dy})^2} dy \]

OpenStudy (nathanjhw):

The limits are 0 and 2 right?

OpenStudy (freckles):

yep

OpenStudy (nathanjhw):

So does the equation become: integral (2,0) (2piy(sqrt(1+2x/y^2)^2)dy

OpenStudy (nathanjhw):

@freckles

OpenStudy (freckles):

how did you get dx/dy to be a function of x and y?

OpenStudy (freckles):

dx/dy should just be a function of y

OpenStudy (freckles):

\[x=\frac{1}{2} y^2 \\ \frac{dx}{dy}= ? \text{ use power rule } \]

OpenStudy (nathanjhw):

1/y?

OpenStudy (freckles):

2-1=1 not -1

OpenStudy (freckles):

\[\frac{dx}{dy}=\frac{1}{2} \cdot 2 y^{2-1}=\frac{2}{2}y^{1}=1y^1=y^1 =y\]

OpenStudy (nathanjhw):

So if dx/dy =y. Then the equation becomes integral (0,2) (2piy(sqrt((1+(sqrt2x))^2))dy

OpenStudy (freckles):

how are you getting that

OpenStudy (freckles):

replace dx/dy with y

OpenStudy (freckles):

\[\int\limits\limits_{c}^{d}2 \pi y \sqrt{1+(\frac{dx}{dy})^2} dy \\ \int\limits_0^2 2 \pi y \sqrt{1+(y)^2} dy \]

OpenStudy (freckles):

you can drop the ( ) around the y since y is just one term

OpenStudy (nathanjhw):

When I integrated that I got 4 π y sqrt(1+y^2)

OpenStudy (freckles):

looks like you just multiplied the integral by 2

OpenStudy (freckles):

that isn't integration

OpenStudy (freckles):

you might want to try a substitution before the actual integration

OpenStudy (freckles):

let u=1+y^2 du/dy=2y

OpenStudy (nathanjhw):

I made an error. I did dx instead of dy. I got 21.322 when I tried again.

OpenStudy (freckles):

sounds much better

OpenStudy (nathanjhw):

Thank you very much!